Malik canoed against the current for 2 hours and then with the current for 1 hour. Julio traveled against the current for 2.5 hours and with the current for 1.5 hours. If they traveled a total of 9.5 miles against the current and 20.5 miles with the current and the current is going 3 miles per hour how fast do Malik and Julio travel in still water?
Answer by mananth(16946)

Malik canoed against the current for 2 hours and then with the current for 1 hour. Julio traveled against the current for 2.5 hours and with the current for 1.5 hours. If they traveled a total of 9.5 miles against the current and 20.5 miles with the current and the current is going 3 miles per hour how fast do Malik and Julio travel in still water?
Let Malik's rate be x mph
and Julio's rate be y mph
rate of current = 2 mph
Malik upstream distance = (x-3)*2
Jukio's upstream distance = (y-3)*2.5
2(x-3)+2.5(y-3)= 9.5
2x-6+2.5y-7.5= 9.5
2x+2.5y -13.5=9.5
2x+2.5y=23
multiply by 2
4x+5y=46................(1)

Similarly
down stream
1(x+3)+1.5(y+3)=20.5
x+3 +1.5y+4.5 =20.5
x+1.5y=13
multiply by -4
-4x-6y=--54...........(2)
Add (1) & (2)
-y= -8
y= 8


4x+5y=46
plug y
we get 4x+40=46
x=1.5